40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Fri Oct 21 00:19:36 2011, Dmass, DailyProgress, Laser, Planex Laser Work 
    Reply  Sun Oct 23 18:16:51 2011, rana, DailyProgress, Laser, Planex Laser Work 
    Reply  Tue Oct 25 20:01:01 2011, Dmass, DailyProgress, Laser, Planex Laser Work [editing] CoherenceTime.pdfCharTimeSers10_21_2011.pdfZeroCrossings.pdfHistAll.pdfBeatFreqNoiseAll.pdf
    Reply  Wed Oct 26 03:14:10 2011, Dmass, DailyProgress, Laser, Planex Laser Work 
       Reply  Wed Oct 26 23:01:21 2011, Dmass, DailyProgress, Laser, Planex Laser Work NoiseFilt.pdfLinewidthCompareFilt.pdf
          Reply  Fri Oct 28 16:49:24 2011, Dmass, DailyProgress, Laser, Planex Laser Work 
Message ID: 313     Entry time: Fri Oct 21 00:19:36 2011     Reply to this: 317   320   321
Author: Dmass 
Type: DailyProgress 
Category: Laser 
Subject: Planex Laser Work 

Spent some time this week working towards building a PLL and characterizing the PLANEX diodes. Blow By Blow Summary:

Diode Mounting:

  1. Got some extra Thorlabs LM14S2 (web page) butterfly diode mounts
  2. Checked PLANEX datasheet (elog:134) for its pinout
  3. The PLANEX is TYPE 2 in LM14S2 nomenclature, this chooses which adapter board insert goes inside the LM14S2
  4. Grounded myself with some wire, then removed the PLANEX diodes from the package
  5. Cleaned PLANEX and LM14S2 with methanol
  6. Spread a thin layer of Artic Silver Thermal Paste (web page) on the LM14S2
  7. Placed PLANEX in LM14S2, wiggled it around to spread the silver paste

Cabling:

  1. Only had 1 working set of cables for the ITC510 (web page)
  2. Checked pinout carefully
  3. Pinout is from manuals, for the current cable its 1:1 D
  4. Decapitated an D15 cable to make the D15 -> D9 TEC cable
  5. For the TEC its a little trickier, the cable I made should work for any type of diode in the LM14S2/ITC510 interface:
D15 Male Pin # D-9 female             Connection
2,5,6,7 4 TEC +
9,13,14,15 5 TEC - / / TEC lockout
3 3 Thermistor - (Gnd)
4 2 Thermistor +
1 1 Status LED +
10 7 AD590- (not in PLANEX)
8 8 Status LED - (Gnd)
11 9 AD590+ (

Checking the PLANEX:

From the datasheets and emails with RIO:

Property Diode 102085 Diode 102068
Rth(Tset) 8110 kOhms = 30 C 8997 kOhms = 27.83 C
Ibias 135 mA 105 mA
Imax 150 mA 125 mA
Pout 13.5 mW 12.5 mW
Lambda 1550.118 1550.127
dLambda/dT* ~27 pm / Kelvin ~25 pm / Kelvin
dLambda/dI 0.19pm / mA 0.31 pm / mA

 * The datasheets (with measurements) they provided showed hysteresis in lambda(T) for each diode - error is ~+/- 50% for dLam/dT

  1. The wavelength difference they measured corresponds to ~0.3 Kelvin of heating on 102085
  2. I turned on the TEC feedback, one was oscillating, one was not
  3. I tuned the PID on each ITC510 via screws on the front of the unit to get oscillations below the visible level on the readout
    1. Might want to do this for real if we end up using these for any length of time via the temp readout on the back
  4. PID levels for both identical:
    1. P: all the way CCW *** I: 1/4 turn CCW from all the way CW  *** D: 1/2 turn CCW from all the way CW
    2. Full range is 5/8 turn for all knobs
  5. Adjusted I_lim manually with screwdriver so the units can't exceed maximum specced current
  6. Adjusted BIAS to zero on both units (this is for the monitor PDs)
  7. Turned on LD's measured power out with Thorlabs power meter:
    1. 102085  ****  Rth = 8.110k  ****  Ibias = 134.9mA  ****  Pmeas = 13.04 mW
    2. 102068  ****  Rth = 8.997k  ****  Ibias = 105.1mA  ****  Pmeas = 12.82 mW
    3. Seems close enough to the specified values!

The Beat Signal:

  1.  Set up the diodes to combine/interfere on a beamsplitter
    1. Made path lengths equal by eye so so they are automatically mode matched from the collimators
    2. Put two steering mirrors between one collimator and beamsplitter for alignment
    3. Blocked one arm
    4. Got beam position near and far field
    5. Blocked other arm
    6. Aligned other beam to overlap with position by eye
    7. Signal seemed to overlap well on a card doing the block 1 / block other test
  2. Aligned onto a thorlabs PDA10CS (web page) diode
  3. Played with current / temperature of 102068
  4. Found beat at 1 MHz using the following settings:
    1. 102068  ****  Rth = 9.298k  ****  Rth_set = 9.398k  **** Ibias = 104.0mA
    2. 102085  ****  Rth = 8.114k  ****  Rth_set = 8.119k  **** Ibias = 134.3mA
  5. Adjusted contrast to 86% with steering mirrors
    1. I forgot to tuned the polarizations, so this got better later
  6. I took some data on the scope and played with it a bit to make sure what I was seeing made sense

 

 

PLOTS TO BE POSTED!!

ELOG V3.1.3-